OpenStudy (anonymous):

Factor completely, then place the factors in the proper location on the grid. x2 - 8x + 16

OpenStudy (anonymous):

=(x-4)(x-4) =(x-4)²

OpenStudy (anonymous):

Can you explain it for me please?

OpenStudy (anonymous):

x²-8x+16 =x²-4x-4x+16 =x(x-4)-4(x-4) =(x-4)(x-4) =(x-4)²

OpenStudy (anonymous):

Thanks! =D

OpenStudy (anonymous):

You're welcome